Project

General

Profile

summary step4 readiness » History » Revision 22

Revision 21 (Nebrensky, Henry, 02 August 2016 17:27) → Revision 22/36 (Hodgson, Paul, 09 August 2016 08:27)

h1.    Summary of Step4 readiness actions list 

 h2. Beamline  

 +Target+ 
 * *Target Alarm handler #1613* "update 22":http://micewww.pp.rl.ac.uk/issues/1613#note-22 (*CAM*) 
 * *Return/replace @target1ctl@* 
 * Updated target manual (*EO/JL*) 
 * *Clean-up of R78 target control crate (DL)* and re-instate facility (possible move to R9) 
 * Beam-bump tuning and re-issue of Target checklists (*MBC*) 

 +Conventional Magnets+ 
 * Restore cooling water 
 ** *review existing water flows and temperature range (MBC)* 
 ** need max current limits for quads in control software? (CAM) 
 * *Calibrate D1 & D2 current reading and update IOC software* (*DL*) 
 * Diagnose failed Danfysik 9000 supply from Q2 (*ISIS*) 
 ** *Have quote for replacement board* 
 * Q1 & Q2 alarm on voltage at zero amps (*CAM*) 
 * Q4 oscillation below 10A and anomalous ramp rate (*DL*) 
 * Earth leakage tests on all magnets (*RHale*) 
 * Q6 leak : too fragile, wouldn't touch it - shifters to watch 
 * Leak sensors and drip trays 
 * *"All Off" controls misbehaviour* 
 * *Water, temperatures and MPS water flows in ALH* (*CAM*) 
 ** Magnets trips should be reported in the ALH (*Pierrick*) 

 +Decay Solenoid+ 
 * *Calibrate readback current reading and update IOC software* (*DL*) 
 * Replace final battery cell (*DL* - _confirmed still outstanding_ ) 
 * Diodes in BBU circuit (*DL*) 
 * Boosted air (*MBC*) 
 ** *Trench sensor _(awaiting read-out)_* 
 ** Compressor sensor 
 ** Backup compressor for DS 
 * Linde PC install: awaiting validation and IOC (*DK*) - _licence issues_ 
 * ALH/archiver/SM corrections #1710 (CAM)? 

 +Other+ 
 * *BLOC training and experience* (*Henry*) 
 * BLOC roster 

 h2. Detectors, Monitoring and Control 

 * MDR7: TOF2: current trip in TOF2_103 
 * -Monitoring Monitoring GUI for the tracker volumes needs to be added in the AppLauncher (*Pierrick*)- (*Pierrick*) 
 * -Issue Issue with the EMR DAQ (restarted the crate) - future?: is there an usable spare? (*Francois*)- (*Francois*) 
 * CDB vs Spreadsheet    +(All)+    -- resolved to remove spreadsheet by end of user run 
 * Alarm Handlers and State Machines: is it ready? 
 * *Confirm ISIS PVs and applications OK in MLCR* (*CAM/Steve*) 

 h2. Cooling Channel 

 * "Commissioning Manager" and planning of that process 
 * RA, Operating Instructions, etc. 
 ** Progress on SS procedure documentation (FC done, Need SS now) 
 * Install FC monitoring in ISIS MCR (*Pierrick/Paolo*) 
 * 1 Hall probe has communication problems @HP-67@ 
 * There needs to be someone on call for the SC magnets. It will be a combination between Josef/Mark 

 h2. Hall and infrastructure 

 * PPS issues #1818 ? 
 * All valid PPS meshing and gates in place 
 * camhall03 inverted 
 * *Shared services/environment* (water, air, computing/networking, UPS ( #1213 ) *in Alarm Handler and in Archiver* (*CAM*) 
 ** Neutron Monitor etc. must alarm on a detected fault as well as on high reading 
 ** Ambient temperature readings and alarms 
 * Post-MDR7: Aircon unit assignment incorrect in monitoring GUI: 2 and 3 are swapped 
 * Aircon unit 4 has some mechanical issue - occasional squealing from fan belt: planned intervention during the shutdown (*John G.*) 

 h2. General Computing 

 * Fix DNS: servers give inconsistent results  
 * Heartbeat vs. Nagios and alarms 
 * Archiver not usable from Remote Monitoring Room (need to squeeze into MLCR for ad-hoc queries) 
 * OnMon/OnRec error messages all need timestamps, to allow correlation (*Durga*) 
 * Data Quality Flags #1383    (*All, Janusz, Pierrick*) 
 * Reco: Data Quality flags #1585    (*Durga, Pierrick*) 
 * @miceraid5@ disk: get in touch with Yordan on self-cleanup (clean with cron/tmpwatch everything untouched for over a month). (*Yordan*) 
 * CDB: do we have all necessary information in the CDB 
 ** Get Target Depth, Target Delay, Total Beam Loss to read out (currently 0 or blank) - _not appropriate for CDB?_ (*Durga/Pierrick*) 
 ** Need something monitoring that this info is live and available in the first place 
 * CDB *viewer* needs to have the cooling channel magnets and absorber status. Action: (*Janusz*) 
 * Implement a search function for the CDB (*Someone*) 
 * CDB corrections backlog from previous user runs  
 * File Compactor:  
 ** Alarm Handler #1236 
 ** Data Quality Flags #1383 (*Henry, Pierrick*) 
 * -Automated RECO mover-: 
 ** -Automated mover and concentrator running- 
 * *Hardware token monitoring and alarms* #1860 
 ** *On-site response* 

 h2. Operations 

 * *Ensure up-to-date documentation and checklists* in MLCR: 
 *** especially [[Control_Room_Operations|Control Room Operations]] which is incomplete and out of date 
 *** need Alarm Handler training (screenshots, etc.?) and File Compactor 
 ** New elements  
 *** Whatever is needed for the SS and FC 
 *** Alarm Handler 
 * Permissive list of MOM-authorisable actions -- in progress 
 * [[Web_Whiteboard|Web Whiteboard]] 
 * [[RunPlans|Settings and Run Plans]] defined 

 |/2=.*Issue*|/2=.*Responsible Party*|/2=.*System*|/2=.*Issue Description*|/2=.*Data Noted*|/2=.*Status*|/2=.*Date Completed*| 
 | 
 |MDR1-DET-05|M.Uchida|Tracker|Sufficient spares for the tracker readout should be commissioned so that they are hot-swappable|2015-01-21|In progress|| 
 | 
 |MDR1-CAM-03|P.Hanlet|EMR|Remote operation of the EMR should be established|2015-01-21|In progress|| 
 | 
 |MDR1-CAM-12|P.Hanlet|RunControl|Consider adding -Online Monitoring- and Online Reconstruction notifications #1802 |2015-01-21|In progress|| 
 | 
 |MDR1-CAM-14|P.Hanlet|Target|Add target optical readouts to Alarm Handler|2015-01-21|Awaiting Matt's fixing PV updating||